OpenStudy (anonymous):

Find an equation in standard form for the hyperbola with vertices at (0, ±10) and asymptotes at y = ±(5/4) x.

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

well, what have you got so far?

OpenStudy (anonymous):

the formula: (x-h)^2/a^2 - (y-k)^2/b^2 = 1 and for asymptotes: y-k = +- (b/a) (x-h) Is that right? if so I am trying to figure out to plug in and solve.

OpenStudy (amistre64):

its mostly right ... we should make the formula more general, since the y parts are not always subtracted. \[\pm\frac{(x-h)^2}{a^2}\mp\frac{(y-k)^2}{b^2}=1\]

OpenStudy (anonymous):

ok so now what?

OpenStudy (amistre64):

also, your asymptotes are actually simplified versions of b/a so there might be a common factor to adjust with so that: \[y-k=\pm\frac{bn}{an}(x-h)\]

OpenStudy (amistre64):

well, what do h and k represent to us?

OpenStudy (anonymous):

the center i believe

OpenStudy (amistre64):

good, center is the point (h,k) and what would our center be for this?

OpenStudy (anonymous):

at the origin

OpenStudy (amistre64):

so 0,0 ... we should also determine if the x and y parts in our formula are + or - how do we know which part is subtracted from the other?

OpenStudy (anonymous):

whether the hyperbola is horizontal or vertical? I think

OpenStudy (amistre64):

youre doing good ... what would you say is the correct orientation for this one then?

OpenStudy (anonymous):

 I think is what it would look like, so vertical?

OpenStudy (amistre64):

vertical, or rather .. opens in the y direction, to y + and x is - \[\frac{y^2}{(bn)^2}-\frac{x^2}{(an)^2}=1\] \[\frac{bn}{an}=\frac54\] so, we only need to determine bn, an ... what are your thoughts for it?

OpenStudy (amistre64):